OpenStudy (fanduekisses):

The front face of a Nordic house is triangular. The surface area of the face is x^2 + 3x + 10 where x is the base of the triangle. What is the height of the triangle in terms of x?

OpenStudy (fanduekisses):

1/2 * x * h= x^2 + 3x + 10 right? but then idk what to do :(

OpenStudy (anonymous):

dude enough talking ok i have a plan

OpenStudy (fanduekisses):

?

OpenStudy (anonymous):

k what grade are you in

OpenStudy (fanduekisses):

11th

OpenStudy (fanduekisses):

this is algebra 2

OpenStudy (anonymous):

i am only in 9th grade but i know it give me a few seconds

OpenStudy (fanduekisses):

well that's awesome :D yeah there are a couple of 9th graders in my class lol

OpenStudy (anonymous):

h = 2x + 6 + 20/x t = V/(pi r^2) t = d/S1+d/S2 S = 2/(1/S1+1/S2)

OpenStudy (anonymous):

there is your answer

OpenStudy (fanduekisses):

-.- i know you just copied and pasted that

OpenStudy (fanduekisses):

off yahoo answers lol cos the other two answers are from a different question...lol

OpenStudy (anonymous):

or 1/2*x * h = x^2 + 3x + 10 so the height h =2( x + 3 + 10/x) thick of oil on the surface h V = 4/3TTr^2h h = 3V/44TTr^2

OpenStudy (fanduekisses):

and I already knew the answer, just wanted to know how to solve it -.-

OpenStudy (anonymous):

well there is the work

OpenStudy (anonymous):

give me my medal please

OpenStudy (fanduekisses):

errg ill just give it bc i'm feeling nice but not bc i found your answer helpful

OpenStudy (anonymous):

thanks